OpenStudy (hanna1):

plz help will medal and fan!! Solve for Y 4x-5y=9

HanAkoSolo (jamierox4ev3r):

First, you want to get "y" on one side of the equation. Any ideas on how to do this? @hanna1

OpenStudy (hanna1):

not with this problem

HanAkoSolo (jamierox4ev3r):

okay. First, we want to subtract 4x from both sides, like this: 4x-5y=9 -4x -4x -------- -5y = -4x+9

OpenStudy (hanna1):

okay?

HanAkoSolo (jamierox4ev3r):

Does that make sense so far?

OpenStudy (hanna1):

nope cuz u can subtract 4x from 9

HanAkoSolo (jamierox4ev3r):

no, you add the 4x to the 9...it's basically moving 4x to the other side of the equation

OpenStudy (hanna1):

oh okay

HanAkoSolo (jamierox4ev3r):

Now, do you have any ideas on what to do next? (Hint: we have only one more step!)

OpenStudy (hanna1):

nope

HanAkoSolo (jamierox4ev3r):

Okay. You want to divide everything by -5 to isolate y.

OpenStudy (hanna1):

okay

HanAkoSolo (jamierox4ev3r):

It would look like this: -5y = -4x+9 /-5 /-5 /-5 ------------ ? Try to see what you would get. I'll help you if needed

OpenStudy (hanna1):

thx

HanAkoSolo (jamierox4ev3r):

no problem. What did you get?

OpenStudy (hanna1):

y=4/5x-9/5

HanAkoSolo (jamierox4ev3r):

correct :)

OpenStudy (hanna1):

thx so much<3

HanAkoSolo (jamierox4ev3r):

no problem :)